# Diffraction Aided Wireless Positioning

### v1

@@ Abstract
Wireless positioning in Non-Line-of-Sight (NLOS) scenarios is highly challenging due to multipath, which leads to deterioration in the positioning estimate. This study reexamines electromagnetic field principles and applies them to wireless positioning, resulting in new techniques that enhance positioning accuracy in NLOS scenarios. Further, we use the proposed method to analyze a public safety scenario where it is essential to determine the position of at-risk individuals within buildings, emphasizing improving the Z-axis position estimate. Our analysis uses the Geometrical Theory of Diffraction (GTD) to provide important signal propagation insights and develop a new NLOS path model. Next, we use Fisher information to derive necessary and sufficient conditions for 3D positioning using our proposed positioning technique and finally to lower bound the possible 3D and z-axis positioning performance. On applying this positioning technique in a public safety scenario, we show that it is possible to greatly improve both 3D and Z-axis positioning performance by directly estimating NLOS path lengths.
@@
Geometric theory of Diffraction, time-of-flight, wireless positioning, emergency networks, Z-axis positioning, 3D localization, outdoor-to-indoor propagation.
@@ I Introduction
In wireless systems, Time-of-Flight (TOF) based techniques are ubiquitous for estimating position. Therefore this technique sees widespread deployment in various positioning technologies, including GNSS systems [2], WiFi networks, 4G and 5G networks [3], etc. In its most general form, we have several anchors with position knowledge and a node at an unknown position, which is to be estimated. TOF measurements are conducted between the anchors and the node to measure the distance between the anchor and the node. We can obtain 2D or 3D position estimates for the node using these measurements. This system works well when Line-of-Sight (LOS) conditions exist when the signal propagation time corresponds to the Euclidean distance between the anchor and the node. However, in Non-Line-Of-Sight (NLOS) scenarios, we face numerous distinctive challenges. The wireless signal propagates along paths in 3D space and these are known as Multipath components (MPC) of the signal. The presence of non-resolvable MPCs leads to fluctuations in the received signal strength, and in the case of resolvable MPCs, we can improve the position estimate under certain conditions [4]. In NLOS scenarios where the direct path between the anchor and node does not exist, the NLOS paths are longer than the direct path. TOF-based ranging measurements conducted in this scenario will estimate a path length longer than the direct path, and the excess path length is termed NLOS bias. NLOS bias leads to degraded positioning performance and is an important challenge that needs to be resolved, especially for indoor positioning.
@@
Previous work has tackled the problem of NLOS bias using two broad techniques - modeling it either as (a) a stochastic quantity or, (b) as an unknown deterministic quantity that can be estimated under certain conditions. O’Lone et al. [5] showed mathematically that the NLOS bias due to single-bounce reflections follows the exponential distribution if modeled as a random variable. This approach provides a tractable way to quantify a priori information associated with the NLOS paths. Shen et al. [4] showed that in the absence of a priori information about the NLOS bias distribution for the NLOS paths, we can discard the NLOS measurements corresponding to these NLOS paths. However, Qi et al. [6] show that the position estimate improves by incorporating a priori information about the NLOS bias. This led to development of algorithms that are robust to NLOS bias [7, 8, 9, 10].
@@
Moving to the second approach, where NLOS bias is modeled as a deterministic unknown quantity, Witrisal et al. [11] developed a framework to incorporate information from floor map plans into TOF-based measurements to improve indoor positioning estimates. Leitinger et al. [12] derived the Cramér-Rao Lower Bound (CRLB) to lower bound the achievable performance for this positioning methodology for the indoor positioning scenario. Mendrzik et al. [13] assumed NLOS paths were generated due to single bounce specular reflections and proposed to use an antenna array in conjunction with TOF measurements to estimate the point of reflection. This enabled the direct estimation of NLOS path lengths, leading to improved position estimates. In the absence of a priori map information and the ability to directly estimate the point of incidence for the NLOS paths, several other works have used collaborative techniques [14] or SLAM-based techniques [15, 16] in conjunction with various measurements corresponding to different MPCs to localize a node within an unknown environment and simultaneously developing a map of the environment. Nevertheless, the problem of NLOS bias remains a significant challenge in modern wireless positioning systems.
@@
A critical use case in modern wireless networks is positioning for public safety scenarios, which poses numerous distinctive challenges. Wireless networks can become highly congested during emergencies such as mass shootings or firefighting incidents, resulting in suboptimal coverage and adversely impacting emergency response efforts. From the industrial perspective, a solution to address this challenge has been devised in the United States through a public-private partnership between the federal government and AT&T and is called FirstNet [17]. To counter network congestion, FirstNet proposed the use of a dedicated band - LTE Band 14 exclusively for public safety usage. Further, the Z-axis position of the UE is measured using a barometer and reported over the dedicated band for positioning purposes. However, this solution relies on fixed cellular base stations, making them susceptible to potential damage, especially during emergencies such as fires, which can result in loss of coverage due to power loss.
@@
Therefore, there is a pressing need for a swiftly deployable private wireless network tailored to meet the exclusive requirements of public safety for both communication and positioning. Such a network would enhance indoor coverage and reliability, thereby contributing to developing effective emergency response strategies. Prior art has proposed a system that involves the deployment of mobile UAVs equipped with positional awareness, denoted as ‘anchors.’ These mobile anchors establish wireless connections with User Equipment (UE), called ‘nodes’, located within a building [3, 18, 19, 1]. Previous analysis of this system demonstrates how this system can enhance indoor coverage by leveraging UAV mobility, as discussed in [19], while addressing indoor positioning requirements using 5G technology, as elaborated in [18]. However, a key challenge for indoor scenarios lies in further refining the accuracy of the Z-axis position estimation, especially as quickly navigating between floors poses challenges for emergency responders. The primary source of the position error is due to the presence of NLOS bias in the TOF measurements.
In the absence of LOS, NLOS paths have been modeled assuming specular reflections as the primary propagation mechanism. Hence, using Snell’s laws, NLOS anchors can be reflected across planar reflecting surfaces (such us walls), leading to virtual anchors. This is analogous to increasing the number of anchors and improving the anchor geometry, thus improving positioning performance. However, there are other propagation scenarios where diffraction from edges might be more common. For example, in Outdoor-to-indoor (O2I) scenarios, diffraction from window edges may actually be more common. In this paper, we investigate diffraction as a propagation mechanism, and our analysis involves Electromagnetic (EM) Field Theory and Fisher Information-based 3D-positioning analysis. Our main contributions are outlined below.
  • Diffraction inspired Path Length Model: By modeling diffraction using electromagnetic field theory, we draw new connections to the field of positioning using wireless signals which results in a path length model.
  • New Positioning Technique for NLOS scenarios: Using the proposed path length model, we create a new positioning technique for NLOS scenarios based on TOF signal measurements. We analyze this positioning technique using a mathematical framework based on Fisher information. Our analysis leads to the derivation of the necessary and sufficient conditions for 3D position estimation and the determination of the CRLB for the 3D position estimation for this technique.
  • Application to a public safety scenario: We apply the proposed positioning technique to a public safety scenario where there is a need to estimate the 3D position of emergency responders within buildings, emphasizing improving the Z-axis position estimate. For this scenario, we present insights about the signal propagation and various system implementation details to improve the indoor positioning estimate. Our central insight is that the first arriving path at indoor locations is due to diffraction from window edges. Further, we favorably influence the O2I signal propagation by controlling the electric field polarization at the transmitter, orienting the receiver side antenna to point towards windows to boost the window diffraction MPCs, and presenting various other system insights.
@@
Our analysis starts by presenting preliminaries from electromagnetic field theory to model diffraction from edges with the goal to obtain path length and electric field magnitude corresponding to paths resulting from diffraction. We use the results derived in this section to develop a simplified building model with diffraction as the primary propagation mechanism, based on which we develop a TOF based positioning technique. Consequently, we develop a Fisher information framework to analyze the proposed positioning technique and conclude the paper by presenting a public safety scenario where we show improvement in both the 3D and Z-axis positioning performance.
@@ II Asymptotic Techniques To Model Diffraction
In this section, we present some preliminaries from electromagnetic field theory. Using the Geometrical Theory of Diffraction (GTD) [20, 21, 22], we first present a canonical example known as 3D edge diffraction for a half plane to model diffraction from edges. Rather than numerically evaluating Maxwell’s equations, classical Geometrical Optics (GO) models refraction and reflection phenomena using ray propagation governed by simple geometrical rules. GTD extends GO, enabling the modeling of diffraction phenomena through ray propagation with similar straightforward geometrical principles. Another asymptotic technique to model diffraction is the Knife Edge Diffraction (KED) which in contrast to GTD does not include the electric field polarization aspects. Previously, GTD has been extensively adopted to model signal attenuation due to propagation in different scenarios concerning edges. In [23],[24], the diffraction losses around buildings were modeled using GTD, and the predictions were validated through measurement campaigns. Pallaprolu et al. [25, 26] use GTD to model edge diffraction for edge imaging applications.
@@
In this section, we aim to derive closed form expressions for both the the path length and the electric field strength of the signal propagation due to diffraction from an edge. Consider Fig. 1(a); we have a perfectly conducting half-plane at $Y=0$ whose edge is a distance $z_{e}$ from the origin and parallel to the X-axis. Note that the edge itself is considered to be infinite in length; however, in practice, a finite length corresponding to several wavelengths is sufficient. The source is at point $A$, the observation is at point $B$, and both these points are located on different sides of the half-plane. The path the diffraction electric field follows is along ray $AQ_{e}$ and then $Q_{e}B$ such that the point $Q_{e}$ lies on the diffracting edge. The location of the point $Q_{e}$ on the edge is such that the distance $AQ_{e}B$ is minimized. This is known as Fermat’s principle of least time [27]. Therefore, the location of the diffraction point $Q_{e}$ is ultimately dependent on both the relative position of the source and observation point with respect to the diffracting edge. Fermat’s principle can also be shown equivalent to the diffraction law [21, 22], and we use this to derive a closed-form expression of the path length $AQ_{e}B$. The diffraction law predicts that the incident ray on an edge leads to a cone of diffracted rays. This cone has been observed under certain conditions [28] and is referred to as the Keller cone. Pallaprolu et al. [26, 25] demonstrated imaging of edges by observing conic sections formed by the intersection of the Keller cones and the receive side planar antenna arrays. In terms of the diffraction law, the location of the tip of the Keller cone $Q_{e}$ is dependent on the relative location of the source point A and observation point B with respect to the edge. Next, we choose the diffracted ray on this Keller cone that passes through the observation point $B$. Hence, $AQ_{e}B$ represents the path followed by the diffracted field to reach point $B$.
Now, we express the diffraction law in more mathematical terms by defining the edge vector $\hat{\bm{e}}$ to point in the tangential direction to the edge. To resolve the ambiguity of the two possible tangential directions, we define two other vectors – $\hat{n}_{0}$ as normal to the half plane in the direction of the source point $A$ and $\hat{t}_{0}$ as the normal to the edge vector and pointing away from the edge along the surface of the half-plane. These three orthonormal vectors are shown in Fig. 1(a) and satisfy
$\hat{t}_{0}=\hat{n}_{0}\times\hat{e}.$ (1)
(a) Edge diffraction using GTD for a half plane
(b) Simplified model of the building
@@
The law of diffraction states that the angle $\gamma_{0}$ between the incident ray $\hat{s}^{{}^{\prime}}$ and the edge $\hat{e}$ is the same angle between the diffracted ray $\hat{s}$ and the edge $\hat{e}$. This is mathematically expressed as
@@
Here $\hat{s}^{\prime}$, $\hat{s}$ and $\hat{e}$ are the unit vectors along the incident ray, diffracted ray, and edge respectively and $\overrightarrow{A}\cdot\overrightarrow{B}$ denotes the dot product between vectors $\overrightarrow{A}$ and $\overrightarrow{B}$.

### v2

# Diffraction-Aided Wireless Positioning

@@ Abstract
Wireless positioning in Non-Line-of-Sight (NLoS) scenarios presents significant challenges due to multipath effects that lead to biased measurements and reduced positioning accuracy. This paper revisits electromagnetic field theory related to diffraction and in the context of wireless positioning and proposes a novel positioning technique that greatly improves accuracy in NLoS environments dominated by diffraction. The method is applied to a critical public safety use case: precisely locating at-risk individuals within buildings, with a particular focus on improving 3D positioning and z-axis accuracy. By leveraging the Geometrical Theory of Diffraction (GTD), the approach introduces an innovative NLoS path length model and a new NLOS positioning technique. Using Fisher information analysis, we establish the conditions required for 3D positioning and derive lower bounds on positioning performance for both 3D and z-axis estimates for the proposed NLOS positioning technique. Additionally, we propose an algorithmic implementation of the proposed NLoS positioning method using non-linear least squares estimation, which we term D-NLS. The positioning performance of our proposed NLOs positioning technique is validated using an extensive ray-tracing simulation. The numerical results highlight the superiority of our approach in outdoor-to-indoor environments, which directly estimates NLoS path lengths and delivers significant performance enhancements over existing methods for both 3D and z-axis positioning scenarios.
@@
Geometric theory of Diffraction, Time-of-Flight, wireless positioning, z-axis positioning, 3D localization, outdoor-to-indoor propagation, NLoS positioning
@@ I Introduction
In wireless systems, Time-of-Flight (ToF) measurements are a widely used geometric-based positioning method, commonly employed for position estimation. As a result, this approach is extensively implemented in various positioning technologies, such as GNSS systems [2], WiFi networks, and 4G and 5G networks [3], among others. In its most general form, we have several anchors with position knowledge and a node at an unknown position, which is to be estimated. ToF measurements are conducted between the anchors and the node to measure the distance between the anchor and the node. We can obtain 2D or 3D position estimates for the node using these measurements. This system works well when Line-of-Sight (LoS) conditions exist when the signal propagation time corresponds to the Euclidean distance between the anchor and the node. However, in Non-Line-Of-Sight (NLoS) scenarios, we face numerous distinct challenges. The wireless signal propagates along paths in 3D space and these are known as Multipath components (MPC) of the signal. The presence of non-resolvable MPCs leads to fluctuations in the received signal strength, and in the case of resolvable MPCs, we can improve the position estimate under certain conditions [4]. In NLoS scenarios where the direct path between the anchor and node does not exist, the NLoS paths are longer than the direct path. ToF-based ranging measurements conducted in this scenario will estimate a path length longer than the direct path, and the excess path length is termed NLoS bias. NLoS bias leads to degraded positioning performance and is an important challenge that needs to be resolved, especially for indoor positioning.
@@
Previous work has tackled the problem of NLoS bias using two broad techniques - modeling it either as (a) a stochastic quantity or, (b) as an unknown deterministic quantity that can be estimated under certain conditions. O’Lone et al. [5] showed mathematically that the NLoS bias due to single-bounce reflections follows the exponential distribution if modeled as a random variable. This approach provides a tractable way to quantify a priori information associated with the NLoS paths. Shen et al. [4] showed that in the absence of a priori information about the NLoS bias distribution for the NLoS paths, we can discard the NLoS measurements corresponding to these NLoS paths. However, Qi et al. [6] show that the position estimate improves by incorporating a priori information about the NLoS bias. This led to the development of algorithms that are robust to NLoS bias [7, 8, 9, 10, 11].
@@
Moving to the second approach, where the NLoS bias is modeled as a deterministic unknown quantity, Witrisal et al. [12] developed a framework to incorporate information from floor map plans into ToF-based measurements to improve indoor positioning estimates. Leitinger et al. [13] derived the Cramér-Rao Lower Bound (CRLB) to lower bound the achievable performance for this positioning methodology for the indoor positioning scenario. Mendrzik et al. [14] assumed that the NLoS paths were generated due to single bounce reflections and proposed to use an antenna array in conjunction with ToF measurements to estimate the point of reflection. This enabled direct estimation of NLoS path lengths, leading to improved position estimates. In the absence of a priori map information and the ability to directly estimate the incidence points for the NLoS paths, several studies have utilized collaborative techniques [15] or proposed SLAM-based approaches [16, 17, 18, 19, 20]. SLAM-based techniques, in particular, leverage the reflection model to simultaneously localize the user and create a map of the unknown environment. This work focuses primarily on the localization aspect, and these SLAM-based approaches could potentially benefit by being extended to include our proposed diffraction-based positioning technique.
@@
A critical use case in modern wireless networks is positioning for public safety scenarios, which poses numerous distinctive challenges. Wireless networks can become highly congested during emergencies such as mass shootings or firefighting incidents, resulting in suboptimal coverage and adversely impacting emergency response efforts. From the industrial perspective, a solution to address this challenge has been devised in the United States through a public-private partnership between the federal government and AT&T and is called FirstNet [21]. To counter network congestion, FirstNet proposed the use of a dedicated band - LTE Band 14 exclusively for public safety usage. In addition, the z-axis position of the UE is measured using a barometer and reported over the dedicated band for positioning purposes. However, this solution relies on fixed cellular base stations, making them susceptible to potential damage, especially during emergencies such as fires, which can result in loss of coverage due to power loss.
@@
Therefore, there is a pressing need for a swiftly deployable private wireless network tailored to meet the exclusive requirements of public safety for both communication and positioning. Such a network would improve indoor coverage and reliability, thus contributing to the development of effective emergency response strategies. Prior art has proposed a system that involves the deployment of mobile UAVs equipped with positional awareness, denoted as ‘anchors.’ These mobile anchors establish wireless connections with User Equipment (UE), called ‘nodes’, located within a building [3, 22, 23, 1, 24]. Previous analysis of this system demonstrates how this system can improve indoor coverage by leveraging UAV mobility, as discussed in [23], while addressing indoor positioning requirements using 5G technology, as elaborated in [22]. However, a key challenge for indoor scenarios lies in further refinement of the accuracy of the z-axis position estimation, especially as quickly navigating between floors poses challenges for emergency responders. The primary source of the position error is due to the presence of NLoS bias in the ToF measurements.
To summarize, when the nature and number of interactions with objects along the NLOS propagation path of a wireless signal are unknown, we model this path using the Euclidean distance with an additive non-zero stochastic NLOS bias term [5]. Alternatively, we can use Snell’s laws to reflect anchors across planar surfaces, leading to virtual anchors. Now, the NLOS signal propagation path can be modeled using the Euclidean distance from the virtual anchor, and this is analogous to increasing the number of anchors. However, there are other propagation scenarios in which diffraction from edges might be more common. For example, in Outdoor-to-indoor (O2I) scenarios, diffraction from window edges may actually be more common. Prior works investigating diffraction as a propagation mechanism have been motivated by the need to explain the non-zero electric field at the receiver when the transmitter and receiver have an obstruction in between them. In the absence of a transmitted path and reflections from other obstacles in the environment, a numerical technique to model wireless signal propagation called geometrical optics (GO) incorrectly predicts a zero electric field at the receiver [25, 26]. In this context, diffraction has been studied to model path losses for signal propagation in building scenarios [27, 28] and for edge imaging applications [29, 30]. In our work, we argue that diffraction from edges plays a dominant role in signal propagation in many scenarios, and thus positioning estimation performance can be improved in NLOS conditions by modeling diffraction in the positioning technique. Specifically, we model the length of the path followed by the diffraction field and incorporate this model in a non-linear least squares positioning technique. We improve on the conference version of this paper [1] with a more in-depth inquiry that includes analysis of the diffraction electric field, Fisher information based analysis of the positioning technique and a very comprehensive raytracing simulation to validate our findings. Our main contributions are as follows:
  • Diffraction-inspired path length model: When an obstacle with an edge lies between the transmitter and receiver, it is plausible that the diffraction field propagates from the transmitter to the edge and then to the receiver. However, the precise location of the diffraction point along the edge is not immediately obvious. By applying electromagnetic principles, we derive a closed-form expression to determine the diffraction point’s location on the edge. This serves as the fundamental connection between modeling diffraction and wireless positioning and leads to a new diffraction-inspired NLOS path model for positioning.
  • O2I Signal Propagation Insights: We argue that diffraction from window edges plays a key role in outdoor-to-indoor (O2I) signal propagation from a transmitter located outside a building to an indoor receiver. To investigate this further we develop a simplified model of a building. Applying the Geometrical Theory of Diffraction to this building model, we predict the presence of two diffraction multipath components (MPC’s) resulting from diffraction from the upper and lower edges of the building windows. Further, by controlling the transmit side polarization we can control the relative power in the two diffraction MPCs and this leads to a single dominating diffraction-based MPC. Lastly, we also predict that this dominating diffraction MPC in the absence of a direct Euclidean path typically is the first arriving path.
  • New NLOS Positioning Technique: Building on the above results and insights, we propose a new NLOS positioning technique which is applicable to an important public safety scenario where there is a need to improve the indoor positioning performance. This public safety positioning system consists of multiple outdoor UAV anchors and indoor UEs (nodes) that need to be localized. First, we modify the proposed general diffraction path length model to make it more suitable to the O2I signal propagation. This modification improves on the conference version of this paper [1] by removing the need to have a priori knowledge of the exact positions of the diffracting edges. Next, we model the range measurements associated with the dominating diffraction MPC using the modified diffraction path length model and leverage Fisher information to derive the necessary and sufficient conditions for estimating the 3D position of the indoor node. The Fisher information can also be used to develop the CRLB to lower bound the positioning performance. We also derive a diffraction-based non-linear-least squares positioning algorithm (D-NLS) that relies on the proposed diffraction path length model. Finally, we validate all the developed results using a comprehensive raytracing simulation that models reflections, diffractions and transmissions for this O2I scenario using a popular wireless modeling software [31]. We demonstrate that we can estimate the delay associated with the diffraction paths using the first arriving path principle which leads to greatly improved 3D positioning performance using the D-NLS algorithm. Note that although the diffraction-based path isn’t always the first arriving path, it is very often the first-arriving path. Also, when it isn’t the first-arriving path, the difference in path length between the first-arriving path and the diffraction-based path is small. This approach is compared both to prior art positioning techniques for NLOS scenarios and also the CRLB.
@@
Our analysis starts by presenting preliminaries from electromagnetic field theory to model diffraction from edges with the goal to obtain the path length and electric field magnitude corresponding to paths resulting from diffraction. We use the results derived in this section to develop a simplified building model with diffraction as the primary propagation mechanism, based on which we develop a ToF based positioning technique. Consequently, we develop a Fisher information framework to analyze the proposed positioning technique and conclude the paper by presenting a public safety scenario where we show improvement in both the 3D and z-axis positioning performance.
@@ II Asymptotic Techniques To Model Diffraction
In this section, we present some preliminaries from electromagnetic field theory. Using the Geometrical Theory of Diffraction (GTD) [32, 26, 25], we first present a canonical example known as 3D edge diffraction for a half plane to model diffraction from edges. Rather than numerically evaluating Maxwell’s equations, classical Geometrical Optics (GO) models refraction and reflection phenomena using ray propagation governed by simple geometrical rules. GTD extends GO, enabling the modeling of diffraction phenomena through ray propagation with similar straightforward geometrical principles. Another asymptotic technique to model diffraction is the Knife Edge Diffraction (KED) which in contrast to GTD does not include the electric field polarization aspects. Previously, GTD has been extensively adopted to model diffraction in different scenarios concerning edges [29, 30, 28, 27, 33].
@@
In this section, we aim to derive closed form expressions for both the path length and the electric field strength of the signal propagation due to diffraction from an edge. Consider Fig. 1(a); we have a perfectly conducting half-plane at $Y=0$ whose edge is a distance $z_{e}$ from the origin and parallel to the X-axis. Note that the edge itself is considered to be infinite in length; however, in practice, a finite length corresponding to several wavelengths is sufficient. The source is at point $\bm{X}_{a}$, the observation is at point $\bm{\alpha}$, and both these points are located on different sides of the half-plane. The path the diffraction electric field follows is along ray $\bm{X_{a}}\bm{Q_{e}}$ and then $\bm{Q_{e}}\bm{\alpha}$ such that the point $\bm{Q_{e}}$ lies on the diffracting edge. The location of the point $\bm{Q_{e}}$ on the edge is such that the geometrical distance $\bm{X_{a}Q_{e}\alpha}$ is minimized. This is known as Fermat’s principle of least time [34]. Therefore, the location of the diffraction point $\bm{Q_{e}}$ is ultimately dependent on both the relative position of the source and observation point with respect to the diffracting edge. Fermat’s principle can also be shown equivalent to the diffraction law [26, 25], and we use this to derive a closed-form expression of the path length specified by the geometrical distance $\bm{X_{a}Q_{e}\alpha}$. The diffraction law predicts that the incident ray on an edge leads to a cone of diffracted rays. This cone has been observed under certain conditions [35] and is referred to as the Keller cone. In terms of the diffraction law, the location of the tip of the Keller cone $\bm{Q_{e}}$ is dependent on the relative location of the source point $\bm{X_{a}}$ and observation point $\bm{\alpha}$ with respect to the edge. Next, we choose the diffracted ray on this Keller cone that passes through the observation point $\bm{\alpha}$. Hence, $\bm{X_{a}Q_{e}\alpha}$ represents the path followed by the diffracted field to reach point $\bm{\alpha}$.
Now, we express the diffraction law in more mathematical terms by defining the edge vector $\bm{\hat{e}}$ to point in the tangential direction to the edge. To resolve the ambiguity of the two possible tangential directions, we define two other vectors – $\bm{\hat{n}_{0}}$ as normal to the half plane in the direction of the source point $\bm{X_{a}}$ and $\bm{\hat{t}_{0}}$ as the normal to the edge vector and pointing away from the edge along the surface of the half-plane. These three orthonormal vectors are shown in Fig. 1(a) and satisfy
$\bm{\hat{t}}_{0}=\bm{\hat{n}}_{0}\times\bm{\hat{e}}.$ (1)
(a) 3D Edge diffraction for a half plane.
(b) Simplified model of the building.
Figure 1: Application of the Geometrical Theory of Diffraction (GTD) (a) Canonical example to model edge diffraction in 3D for a half plane (b) Application of the half plane model to an O2I signal propagation with the transmitter on the UAV-anchor and receiver inside the building.
@@
The law of diffraction states that the angle $\gamma_{0}$ between the incident ray $\bm{\hat{s}^{{}^{\prime}}}$ and the edge $\bm{\hat{e}}$ is the same angle between the diffracted ray $\bm{\hat{s}}$ and the edge $\bm{\hat{e}}$. This is mathematically expressed as
@@
Here $\bm{\hat{s}^{\prime}}$, $\bm{\hat{s}}$ and $\bm{\hat{e}}$ are the unit vectors along the incident ray, the diffracted ray and the edge, respectively, and $\overrightarrow{\bm{A}}\cdot\overrightarrow{\bm{B}}$ denotes the dot product between the vectors $\overrightarrow{\bm{A}}$ and $\overrightarrow{\bm{B}}$.